# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ VII Extension of hardness results for lossy environments
where $\mathcal{I}$ is identity, $\mathcal{E}_{i}$ is an CPTP map representing photon loss, and $\rho_{i}$ is a loss rate for each channel satisfying $\rho_{i}\leq\rho$ for a constant $\rho$. The validity of such modeling for photon loss channel is represented in [15, 3].
[…]
@@
Suppose we have the photon loss model $\mathcal{N}$ with each loss rate $\rho_{i}\leq\rho$ for a constant $\rho$. Then the following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the lossy output probability $p_{\bm{s}}(U,\mathcal{N})$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
We remark that for our noise model, the imprecision blowup grows exponentially with the gate number $m$. Therefore, shallow-depth circuits can be more advantageous in this perspective, since they are likely to have less gate number and thus have small imprecision blowup. For example, the current hardness results are based on $M$ by $N$ submatrices of $M$-dimensional Haar random unitaries, and the implementation of such matrices requires gate number $m=\Omega(N^{\gamma+1})$. This arouses the imprecision blowup at least $2^{O(N^{\gamma+1})}$, which restricts the allowed additive error for the average-case hardness at most $2^{-O(N^{\gamma+1})}$.
VIII Concluding remarks
@@
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in our circuit architecture (i.e., Definition 2) are close enough to those of global Haar random circuits (see, e.g., Ref. [46] and Appendix H). Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [28, 3, 29, 30], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [29]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is the extension of the result in Sec. VII, i.e., finding the classical simulation hardness of noisy Gaussian Boson Sampling with photon loss, or Boson Sampling for more general types of noise. As described in [50, 28, 3], employing the threshold theorem would be a viable choice for this goal, which requires efficient error detection codes using only linear optical elements for these setups. However, to the best of our knowledge, such an error detection code does not yet exist. That is, in contrast to the Boson Sampling case, we cannot post-select the so-called ‘no loss’ event in the Gaussian Boson Sampling case because the input photon number is not fixed in this case. This is also for the Boson Sampling with general types of noise because we cannot post-select no error event by using only photon number measurements. Hence, constructing any of these error detection codes would be a crucial step toward the classical hardness of noisy (Gaussian) Boson Sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with (Gaussian) Boson Sampling. We leave this problem as another open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number $N$. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [1, 3, 27, 28, 30]. This calls for the “dilute” regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [4, 5, 6, 7] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $21\leq\gamma<$ 2 is crucial, similarly to [29] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. IV using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix G, we present an approach that offers useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [46] and Appendix H). If this can be done, we can employ the previous argument [29] that used the global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@
4. We finally remark that if we only use the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ for our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ as defined in Definition 4, the bosonic birthday paradox is not guaranteed because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, for the sake of analytical completeness, we introduced the additional random permutation circuit at the input to guarantee the bosonic birthday paradox and suppress the occurrence of collision outcomes. In fact, from an experimental perspective, this setup can be realized by constructing the local random circuit $\mathcal{H}_{\mathcal{A}}$ and preparing a random input configuration for each experiment. Furthermore, if one can analytically verify that the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ has a permutation invariance property for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, then we can safely leave out this random permutation circuit part, thereby simplifying the average-case circuits and reducing the required circuit depth for the average-case hardness. We leave in Appendix H numerical analysis for the bosonic birthday paradox for our local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, which gives evidence that only the local random circuit ensemble might be sufficient for the average-case hardness result.
@@ Acknowledgements
We thank Bill Fefferman for insightful discussions. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00413957, RS-2024-00438415, NRF-2023R1A2C1006115, and RS-2024-00437191) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2025-RS-2020-II201606). B.G. was supported by the education and training program of the Quantum Information Research Support Center, funded through the National Research Foundation of Korea (NRF) by the Ministry of Science and ICT (MSIT) of the Korean government (No.2021M3H3A1036573). C.O. was supported by the National Research Foundation of Korea (NRF) Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)).
Appendix A Previous foundations on average-case hardness of Boson Sampling
In this appendix, we introduce the existing proof technique used to establish the classical hardness of Boson Sampling, specifically in the context of approximate simulation within a bounded total variation distance error. [1]. The current state-of-the-art proof technique for the hardness of sampling problems like Boson Sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [47]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@ Appendix A Previous foundations on average-case hardness of Boson Sampling
Here, average-case hardness comes into the proof of the classical simulation hardness argument, which means that approximating the ideal output probability of Boson Sampling with high probability over randomly chosen outcomes is #P-hard. More precisely, if the average-case hardness holds up to the imprecision level of the average-case approximation problem, this comes down to the classical simulation hardness of the approximate sampling unless PH collapses, by the complexity-theoretical foundation $\rm{PH}\subseteq\rm{P}^{\#P}$ [48].
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the Boson Sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed due to its translation invariance property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction in terms of circuit instances. Specifically, as the output probability of Boson Sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for Boson Sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [1, 3, 28].
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results and technical developments about the average-case hardness of Boson Sampling [1, 3, 27, 28, 29, 30], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the complete theoretical demonstration of the classical hardness of approximate Boson Sampling.
@@ Appendix B The bosonic birthday paradox for the shallow average-case circuits
Throughout our work, we consider only collision-free outcomes in our average-case hardness arguments instead of all possible outcomes. In this appendix, we examine whether the bosonic birthday paradox [1] holds for our average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$. More specifically, we investigate whether collision-free outcomes are observed with high probability in the so-called “dilute” regime, where the number of modes satisfies $M=\Omega(N^{2})$.
@@
To claim the classical hardness and quantum computational advantage of Boson Sampling based solely on collision-free outcomes, it must be ensured that these collision-free outcomes occupy at least a not-a-very-small portion of the probability weight over all possible outcomes. To understand why, let us denote the probability weight corresponding to the collision-free outcomes by $\varepsilon$. For the probability weight of collision-free outcomes $\varepsilon$, the “adversarial” classical sampler that outputs only collision outcomes is also a good sampler, if the allowed total variation distance error for the classical sampler is larger than $\varepsilon$. Using this adversarial sampler combined with Stockmeyer’s algorithm that multiplicative approximates the output probability of the sampler, the obtained estimates of output probabilities of collision-free outcomes are always “0”, since this sampler does not output collision-free outcomes. Therefore, even if we derive the average-case #P-hardness of estimating output probabilities for collision-free outcomes, we cannot solve this #P-hard problem by Stockmeyer’s reduction if the given sampler is adversarial; thereby, we cannot derive the classical hardness result of sampling (which allows total variation distance error) from the average-case #P-hardness.
@@
Accordingly, to avoid this issue and claim the classical hardness result using only collision-free outcomes, the bosonic birthday paradox is crucial, which suppresses the occurrence of collision outcomes over the experiments. This property holds for global Haar random circuits in the dilute regime $M=\Omega(N^{2})$ [1], but not guaranteed for any other random circuit ensemble, including our circuit ensemble. In fact, collision-free outcomes dominate the outcome space in the regime $M=\Omega(N^{2})$ since the portion of collision-free outcomes over all possible outcomes can be bounded as $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$. Accordingly, one might expect that the bosonic birthday paradox automatically holds for any circuit ensemble. However, even in this regime, the probability weight might be highly concentrated on collision outcomes for some circuit ensembles (i.e., anticoncentration is not guaranteed for arbitrary circuit ensembles), such that the probability weight of the collision-free outcomes is still too small. Hence, it is crucial to show that the bosonic birthday paradox for our circuit ensemble, such that our circuit ensemble outputs collision-free outcomes with high probability.
@@
Based on this understanding, we now argue the bosonic birthday paradox for the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ in the dilute regime $M=\Omega(N^{2})$. We first review a lemma given in Ref. [1, Lemma C.1], which is indeed a crucial ingredient for the bosonic birthday paradox.
Lemma 3 (Unitary Pigeonhole Principle [1]).
Partition a finite set $[K]$ into a “good part” $G$ and “bad part” $B=[K]\setminus G$. Also, let $\hat{U}$ be an arbitrary $K$ by $K$ unitary matrix. Suppose we choose an element $x\in G$ uniformly at random, apply $\hat{U}$ to |$\ket{x}$, and then measure $\hat{U}\ket{x}$ in the standard basis. Then, for measurement outcome $y$, we have $\Pr[y\in B]\leq|B|/|G|$.
Here, let us consider the finite set $[K]$ in the lemma as a set of all possible outcomes of Boson Sampling for $N$ photons and $M$ modes, i.e., the set of possible $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$ (such that, $|K|=\binom{M+N-1}{N}$). Also, we consider $\hat{U}$ in the lemma as a unitary evolution over the Hilbert space of $N$ photons and $M$ modes, characterized by a given unitary circuit of Boson Sampling.
@@
Theorem 8 (Bosonic birthday paradox for shallow random circuits).
@@
We lastly remark that this bosonic birthday paradox is not guaranteed solely by the local random circuit ensemble in Definition 4 itself, because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, the additional random permutation circuit at the input is crucial to guarantee the bosonic birthday paradox and suppress the occurrence of collision outcomes. Nevertheless, if one can prove that the local random circuit ensemble has a permutation-invariance property for a specific circuit architecture, then we can leave out this random permutation circuit part at the input. We leave in Appendix H numerical evidence of the bosonic birthday paradox for solely the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.

### v3

@@ 7 Hardness analyses in lossy environments
where $\mathcal{I}$ is identity, $\mathcal{E}_{i}$ is an CPTP map representing photon loss, and $\rho_{i}$ is a loss rate for each channel satisfying $\rho_{i}\leq\rho$ for a constant $\rho$. The validity of such modeling for photon loss channel is represented in [oszmaniec2018classical, deshpande2022quantum].
To simplify, we assume that we know $\it{a\,priori}$ each error rate $\rho_{i}$ for all $i\in[l]$, and the noise model $\mathcal{N}$ is fixed so that it does not change with random circuit instances. Then we can obtain the hardness of approximating output probabilities of lossy shallow circuits, from our previous hardness proposals. To do so, let $p_{\bm{s}}(C,\mathcal{N})$ be the output probability of an NN-photon outcome $\bm{s}$ from an $M$-mode linear optical circuit $C$ which undergoes loss model $\mathcal{N}$ we set. By post-selecting ‘no loss event’, which can be accomplished by counting the output photon number, the ideal output probability $p_{\bm{s}}(C)$ can be inferred from $p_{\bm{s}}(C,\mathcal{N})$ by
@@
Suppose we have the photon loss model $\mathcal{N}$ with each loss rate $\rho_{i}\leq\rho$ for a constant $\rho$. Then the following problem is $\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the lossy output probability $p_{\bm{s}}(U,\mathcal{N})$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
We remark that for our noise model, the imprecision blowup grows exponentially with the gate number $m$. Therefore, shallow-depth circuits can be more advantageous in this perspective, since they are likely to have less gate number and thus have small imprecision blowup. For example, the current hardness results are based on $M$ by NN submatrices of $M$-dimensional Haar random unitaries, and the implementation of such matrices requires gate number $m=\Omega(N^{\gamma+1})$. This arouses the imprecision blowup at least $2^{O(N^{\gamma+1})}$, which restricts the allowed additive error for the average-case hardness at most $2^{-O(N^{\gamma+1})}$.
8 Concluding remarks
@@
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in our circuit architecture (i.e., Definition 2) are close enough to those of global Haar random circuits (see, e.g., Ref. [go2024exploring] and Appendix I). Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [bouland2022noise, deshpande2022quantum, bouland2023complexity, bouland2024average], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [bouland2023complexity]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix AA as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is the extension of the result in Sec. 7, i.e., finding the classical simulation hardness of noisy Gaussian Boson Sampling with photon loss, or Boson Sampling for more general types of noise. As described in [fujii2016noise, bouland2022noise, deshpande2022quantum], employing the threshold theorem would be a viable choice for this goal, which requires efficient error detection codes using only linear optical elements for these setups. However, to the best of our knowledge, such an error detection code does not yet exist. That is, in contrast to the Boson Sampling case, we cannot post-select the so-called ‘no loss’ event in the Gaussian Boson Sampling case because the input photon number is not fixed in this case. This is also for the Boson Sampling with general types of noise because we cannot post-select no error event by using only photon number measurements. Hence, constructing any of these error detection codes would be a crucial step toward the classical hardness of noisy (Gaussian) Boson Sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with (Gaussian) Boson Sampling. We leave this problem as another open question.
@@
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number NN. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2024average]. This calls for the “dilute" regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [zhong2020quantum, zhong2021phase, madsen2022quantum, deng2023gaussian] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $1\leq\gamma<2$ is crucial, similarly to [bouland2023complexity] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. 4 using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix H, we present an approach that could offer a useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [go2024exploring] and Appendix I). If this can be done, we can employ the previous argument [bouland2023complexity], which used global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
@@
4. We finally remark that if we only use the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ for our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ as defined in Definition 4, the bosonic birthday paradox is not guaranteed because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, for the sake of analytical completeness, we introduced the additional (shallow-depth) random permutation circuit at the input to guarantee the bosonic birthday paradox and suppress the occurrence of collision outcomes. In fact, from an experimental perspective, this setup can be realized by constructing the local random circuit $\mathcal{H}_{\mathcal{A}}$ and preparing a random input configuration. Furthermore, if one can analytically verify that the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ has a permutation invariance property for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, then we can safely leave out this random permutation circuit part, thereby simplifying the average-case circuits and reducing the required circuit depth for the average-case hardness. We leave in Appendix I numerical analysis for the bosonic birthday paradox for our local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, which gives evidence that only the local random circuit ensemble might be sufficient for the average-case hardness result.
@@ Acknowledgements
We thank Bill Fefferman for insightful discussions. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00413957, RS-2024-00438415, and RS-2023-NR076733) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (IITP-2026-RS-2020-II201606 and IITP-2026-RS-2024-00437191). B.G. was supported by the education and training program of the Quantum Information Research Support Center, funded through the National Research Foundation of Korea (NRF) by the Ministry of Science and ICT (MSIT) of the Korean government (No.2021M3H3A1036573). C.O. was supported by the National Research Foundation of Korea Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)) and the Institute of Information & Communications Technology Planning & Evaluation (IITP) Grants funded by the Korea government (MSIT) (No. IITP-2025-RS-2025-02283189 and IITP-2025-RS-2025-02263264). This work was supported by Global Partnership Program of Leading Universities in Quantum Science and Technology (RS-2025-08542968) through the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT(MSIT)).
Appendix A Previous foundations on average-case hardness
In this appendix, we introduce the existing proof technique used to establish the classical hardness of Boson Sampling, specifically in the context of approximate simulation within a bounded total variation distance error. [aaronson2011computational]. The current state-of-the-art proof technique for the hardness of sampling problems like Boson Sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [stockmeyer1985approximation]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@ Appendix A Previous foundations on average-case hardness
Here, average-case hardness comes into the proof of the classical simulation hardness argument, which means that approximating the ideal output probability of Boson Sampling with high probability over randomly chosen outcomes is #P-hard. More precisely, if the average-case hardness holds up to the imprecision level of the average-case approximation problem, this comes down to the classical simulation hardness of the approximate sampling unless PH collapses, by the complexity-theoretical foundation $\rm{PH}\subseteq\rm{P}^{\#P}$ [toda1991pp].
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the Boson Sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed due to its translation invariance property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction in terms of circuit instances. Specifically, as the output probability of Boson Sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for Boson Sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [aaronson2011computational, deshpande2022quantum, bouland2022noise].
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results and technical developments about the average-case hardness of Boson Sampling [aaronson2011computational, deshpande2022quantum, grier2022complexity, bouland2022noise, bouland2023complexity, bouland2024average], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the complete theoretical demonstration of the classical hardness of approximate Boson Sampling.
@@ Appendix B The bosonic birthday paradox for the shallow average-case circuits
Throughout our work, we consider only collision-free outcomes in our average-case hardness arguments instead of all possible outcomes. In this appendix, we examine whether the bosonic birthday paradox [aaronson2011computational] holds for our average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$. More specifically, we investigate whether collision-free outcomes are observed with high probability in the so-called “dilute” regime, where the number of modes satisfies $M=\Omega(N^{2})$.
@@
To claim the classical hardness and quantum computational advantage of Boson Sampling based solely on collision-free outcomes, it must be ensured that these collision-free outcomes occupy at least a not-a-very-small portion of the probability weight over all possible outcomes. To understand why, let us denote the probability weight corresponding to the collision-free outcomes by $\varepsilon$. For the probability weight of collision-free outcomes $\varepsilon$, the “adversarial" classical sampler that outputs only collision outcomes is also a good sampler, if the allowed total variation distance error for the classical sampler is larger than $\varepsilon$. Using this adversarial sampler combined with Stockmeyer’s algorithm that multiplicative approximates the output probability of the sampler, the obtained estimates of output probabilities of collision-free outcomes are always “0", since this sampler does not output collision-free outcomes. Therefore, even if we derive the average-case #P-hardness of estimating output probabilities for collision-free outcomes, we cannot solve this #P-hard problem by Stockmeyer’s reduction if the given sampler is adversarial; thereby, we cannot derive the classical hardness result of sampling (which allows total variation distance error) from the average-case #P-hardness.
@@
Accordingly, to avoid this issue and claim the classical hardness result using only collision-free outcomes, the bosonic birthday paradox is crucial, which suppresses the occurrence of collision outcomes over the experiments. This property holds for global Haar random circuits in the dilute regime $M=\Omega(N^{2})$ [aaronson2011computational], but not guaranteed for any other random circuit ensemble, including our circuit ensemble. In fact, collision-free outcomes dominate the outcome space in the regime $M=\Omega(N^{2})$ since the portion of collision-free outcomes over all possible outcomes can be bounded as $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$. Accordingly, one might expect that the bosonic birthday paradox automatically holds for any circuit ensemble. However, even in this regime, the probability weight might be highly concentrated on collision outcomes for some circuit ensembles (i.e., anticoncentration is not guaranteed for arbitrary circuit ensembles), such that the probability weight of the collision-free outcomes is still too small. Hence, it is crucial to show that the bosonic birthday paradox for our circuit ensemble, such that our circuit ensemble outputs collision-free outcomes with high probability.
@@
Based on this understanding, we now argue the bosonic birthday paradox for the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ in the dilute regime $M=\Omega(N^{2})$. We first review a lemma given in Ref. [aaronson2011computational, Lemma C.1], which is indeed a crucial ingredient for the bosonic birthday paradox.
Lemma 4 (Unitary Pigeonhole Principle [aaronson2011computational]).
[…]
@@
Theorem 8 (Bosonic birthday paradox for the product circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$).
@@
Crucially, the proof of Theorem 8 does not rely on the circuit ensemble $\mathcal{H}_{\mathcal{A}}$; rather, the essential ingredient for the bosonic birthday paradox is the initial random permutation layer $\mathcal{P}$. Therefore, the bosonic birthday paradox is guaranteed in the dilute regime $M=\Omega(N^{2})$ as long as the as long as the input (collision-free) configuration is uniformly random by any means. In practice, this can also be achieved experimentally by directly preparing a uniformly random input configuration. Nevertheless, we chose to include an explicit random permutation circuit $\mathcal{P}$ for mathematical clarity, since we fix the input configuration for Boson Sampling.
We additionally note that this bosonic birthday paradox is not guaranteed solely by the local random circuit ensemble in Definition 4 itself, because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, the additional random permutation circuit at the input is crucial to guarantee the bosonic birthday paradox mathematically. Nevertheless, if one can prove that the local random circuit ensemble has a permutation-invariance property for a specific circuit architecture, then we can leave out this random permutation circuit part at the input. We leave in Appendix I numerical evidence of the bosonic birthday paradox for solely the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.